# Combinatorics of Essential Sets for Positroids

### v2

@@ Abstract
Positroids are families of matroids introduced by Postnikov in the study of non-negative Grassmannians. Postnikov identified several combinatorial objects in bijections with positroids, among which are bounded affine permutations. On the other hand, the notion of essential sets, introduced for permutations by Fulton, was used by Knutson in the study of the special family of interval rank positroids. We generalize Fulton’s essential sets to bounded affine permutations. The bijection of the latter with positroids, allows the study of their relation. From the point of view of positroids, essential sets are maximally dependent cyclic interval. We define connected essential sets and prove that they give a facet description of the positroid polytope, as well as equations defining the positroid variety. We define a subset of essential sets, called core, which contains minimal rank conditions to uniquely recover a positroid. We provide an algorithm to retrieve the positroid satisfying the rank conditions in the core or any compatible rank condition on cyclic intervals.
@@ 1 Introduction
Matroids are combinatorial objects introduced to encode the concept of linear dependence arising from vector spaces, graphs, algebraic field extensions, and various other contexts. Positroids form a particularly well-behaved class of matroids. On the combinatorial side, Postnikov has identified several families of objects in bijection with positroids, such as bounded affine permutations, plabic graphs, Grassmann necklaces, and Le diagrams [13]. The realization space of a positroid inside the totally non-negative Grassmannian, called a positroid cell, is a topological cell homeomorphic to an open ball. Positroid cells fit together to form a CW decomposition of the totally non-negative Grassmannian [14]. Positroids have also been studied in relation to problems coming from quantum physics, as linear images of some positroid cells define a tiling of the amplituhedron, an object encoding the geometry of scattering amplitudes, [2, 6, 11].
Here, we introduce a new combinatorial description of positroids via ranked essential sets and explain how to derive information about the positroid’s structure from them. While many combinatorial objects studied in relation to positroids encode their data implicitly, ranked essential sets provide an intuitive understanding of the positroid’s dependencies. These dependencies are succinctly encoded by relying on the cyclic order of the ground set. Additionally, ranked essential sets offer insights into the geometry of the positroid’s realization space within the Grassmannian and the totally non-negative Grassmannian. As an example, their dimension can be explicitly computed in terms of the ranked essential sets.
@@ 2 Preliminaries
We begin this section by recalling the definitions of matroids and positroids. Specifically, we recall two among the various equivalent definitions of a matroid: one through its maximally independent sets, and the other through its closed sets.
@@
A matroid $\mathcal{M}$ is a pair $(E,\mathcal{B})$ such that $E$ is a finite set, called ground set, and $\mathcal{B}$ is a collection of subsets of $E$, called basis, satisfying:
@@
We will denote the minimal flat containing a subset $S\subseteq E$ by $\langle S\rangle$.
@@
Throughout the paper, without loss of generality, we assume $E=\{1,\dots,n\}=[n]$. The family of cardinality $k$ subsets of $[n]$ is denoted by $\binom{[n]}{k}$. For any $i,j\in\mathbb{Z}$ with $(k-1)n<i\leq kn$ and $(h-1)n<j\leq hn$ for $k,h\in\mathbb{Z}$, we define $[i,j]$ as follows:
  • $\{i\mod n,i+1\mod n,\dots,j\mod n\}$ if ($k=h$ and $i<j$),
  • $\{i\mod n,i+1\mod n,\dots,kn\mod n,kn+1\mod n,\dots,(h-1)n\mod n,(h-1)n+1\mod n,\dots,j\mod n\}$, otherwise.
In other words, $[i,j]$ represents the cyclic interval in $[n]$ from $i\mod n$, to $j\mod n$. For instance, if $n=8$, then $[7,10]=[7,2]=\{7,8,1,2\}$. We use $<$ to denote the normal ordering on $\mathbb{Z}$ and $<_{i}$ to denote the ordering on $[n]$ where $i$ is the smallest element, i.e., $i<_{i}i+1<_{i}\dots<_{i}n<_{i}1<_{i}\dots<_{i}i-1$.
Matroids arise as a generalization of the notion of linear dependency. If a matroid is associated to the dependency behavior of a finite set of vectors in a vector space, it is called realizable. Formally, we define it as follows.
@@
Let $\mathbb{K}$ be a field. A matroid $\mathcal{M}=([n],\mathcal{B})$ is said to be realizable over $\mathbb{K}$ if there exists a full-rank $k\times n$ matrix $M$ with coefficients in $\mathbb{K}$ such that for every $I\in\binom{[n]}{k}$
@@
where $M_{I}$ denotes the square $k\times k$ matrix whose columns are the columns indexed by $I$ in $M$. $M$ is said to be a realizing matrix for $\mathcal{M}$.
A positroid is a realizable matroid over $\mathbb{R}$ with a positivity condition.
@@
A matroid $\mathcal{M}$ is said to be a positroid if it is a realizable matroid over $\mathbb{R}$ such that there exists a realizing matrix $M$ for $\mathcal{M}$ with non-negative maximal minors. In other words, for every $I\in\binom{[n]}{k}$, where $M_{I}$ is as defined in Definition 2.4, we have $\det(M_{I})\geq 0$.
Positroids, along with their combinatorial counterparts, have been subject to extensive research [13]. Among these combinatorial objects bounded affine permutations are of particular interest to us.
@@
Consider a positive integer $n$. A bounded affine permutation of size $n$ is a bijection $\pi:\mathbb{Z}\to\mathbb{Z}$ satisfying the following conditions for all $i,k\in\mathbb{Z}$:
$\pi(i+kn)=\pi(i)+kn\quad\text{and}\quad i\leq\pi(i)\leq i+n.$
@@
$\pi(i)=\min\{j\geq i\mid i\in\langle[i+1,\dots,j]\rangle\},\quad\text{ for every }i\in\mathbb{Z}.$
In particular, if $i$ is a loop of the positroid, then $\pi(i)=i$, and if $i$ is a coloop of the positroid, then $\pi(i)=i+n$. Moreover, information on the rank of a positroid on cyclic intervals can be extracted from its associated bounded affine permutation.
@@
Let $\mathcal{M}=([n],\mathcal{B})$ be a matroid. The matroid polytope is
@@
where $e_{B}$ is the indicator vector of $B$, i.e., $e_{B}=\sum_{i\in B}e_{i}$, with $e_{i}$ vectors of the canonical basis of $\mathbb{R}^{n}$.
@@ 3.1 Essential sets and bounded affine permutations
Consider a $k\times n$ array of squares in the plane. We will always consider the rows modulo $k$, i.e., like a matrix vertically infinitely repeating. The square at coordinates $(i,j)$ corresponds to the square contained in row $i$ and column $j$ of the array. Let $R_{i}$ be the $i$-th row of the array and $A_{(i,j)}$ be the antidiagonal containing the square $(i,j)$. The sub-antidiagonal $\Delta_{(i,j)}$ of the square $(i,j)$ is the set of squares in the array contained in the antidiagonal starting at the square $(i+1,j-1)$ and ending in the first column, with rows considered modulo $k$. Formally, it is defined as:
@@
Let $\pi$ be a bounded affine permutation of size $n$. We associate to it a dotted array as follows: consider an $n\times(n+1)$ array of squares in the plane. For each $i\in[n]$, place a dot inside the square $(i,\pi(i)-(i-1))$, while all other squares remain white. This is well-defined since $i\leq\pi(i)\leq i+n$. We denote by $D(\pi)$ the set of squares in the dotted array of $\pi$ containing a dot, i.e., $D(\pi)=\{(i,\pi(i)-i+1)\mid i\in[n]\}$.
@@
A white square is called a corner if no white square shares its top-right corner. The essential family $\mathcal{E}(\pi)$ of $\pi$ is defined as the set of cyclic intervals $[i,j]$ for each white corner $(i,j-i+1)$ in the diagram of $\pi$. Each element in $\mathcal{E}(\pi)$ is referred to as an essential set. Equivalently:
@@
The positroid corresponding to $\pi$ has rank $3$ and can be represented by the point-line configuration in Figure 2 (right). Note that the essential family captures the sets lying on a line and the parallel elements. However, to distinguish between them, we need to know their rank.
@@
Let $\pi$ be a bounded affine permutation of order $n$. We define the ranked essential family $\mathcal{E}^{r}(\pi)$ of $\pi$ to be the family
$\mathcal{E}^{r}(\pi)=\left\{(r,[i,j])\mid[i,j]\in\mathcal{E}(\pi)\cup\{[1,n]\},\;r=\text{ number of dots in }P_{(i,j-i+1)}\right\}.$
@@
Let $\mathcal{U}_{k,n}$ be the uniform matroid of rank $k$ on $n$ elements. The associated bounded affine permutation is defined by $\pi_{k,n}(i)=i+k$, for every $i\in\mathbb{Z}$. It is easy to see that the set of corners of its diagram is empty, hence $\mathcal{E}^{r}(\pi_{k,n})=\{(k,[1,n])\}$.
@@ 3.2 Essential sets and rank function
The ranked essential family can be interpreted as the rank information of “closed” cyclic interval. Indeed, the essential sets can be constructed from the rank function of the positroid as follows.
@@
Then $\mathcal{E}^{r}=\mathcal{R}\cup\{(k,[1,n])\}$.
@@
Let $\pi$ be the bounded affine permutation associated to the positroid $\mathcal{P}$. Let $(r,[i,j])\in\mathcal{R}$. Since $\operatorname{rk}([i,j])=\operatorname{rk}([i+1,j])$, we have $\pi(i)\leq j$. Similarly, since $\operatorname{rk}([i-1,j])=\operatorname{rk}([i,j])+1$ and $i-1\not\in\langle i,\dots,j\rangle$, we have $\pi(i-1)>j$. It remains to prove that $\pi^{-1}(j)\leq_{i}j$ and $\pi^{-1}(j+1)>_{i}j+1$. The condition $\operatorname{rk}([i,j])=\operatorname{rk}([i,j-1])$ is equivalent to the fact that the number of dots in $P_{(i,j-i+1)}$ and in $P_{(i,j-i)}$ is the same. Then either $\pi(j)=j$ or $\pi(j)>j$. In the latter case, $P_{(i,j-i+1)}$ contains a dot in the last row, which is not part of $P_{(i,j-i)}$. Hence there must be a dot in $P_{(i,j-i)}\setminus P_{(i,j-i+1)}$, i.e. $\pi^{-1}(j)<_{i}j$. It follows that $\pi^{-1}(j)\leq_{i}j$. Finally, $\operatorname{rk}([i,j])=\operatorname{rk}([i,j+1])-1$ implies that the number of dots in $P_{(i,j-i+2)}$ is one more than the number of dots in $P_{(i,j-i+1)}$. Note that $j+1$ is not a loop, hence $\pi(j+1)>j+1$, and $P_{(i,j-i+2)}$ contains a dot in the row $j+1$, which is not in $P_{(i,j-i+1)}$. For the equality to hold there must be no dot in $P_{(i,j-i+1)}\setminus P_{(i,j-i+2)}$, i.e., $\pi^{-1}(j+1)>_{i}j+1$. It follows that $[i,j]$ is an essential set of $\pi$.
Conversely, let $(r,[i,j])\in\mathcal{E}^{r}\setminus\{(k,[1,n])\}$. By definition, $r=\operatorname{rk}([i,j])$ is the number of dots in $P_{(i,j-i+1)}$. Moreover, $\operatorname{rk}([i+1,j])$ is equal to the number of dots in $P_{(i+1,j-i)}$, which is equal to the number of dots in $P_{(i,j-i+1)}$ since there is no dot in the first row of this region by definition of essential set. Therefore, $\operatorname{rk}([i+1,j])=r$. In the same way, since $\pi(i-1)>j$, the number of dots in $P_{(i-1,j-i+2)}$ is equal to one more then the dots in $P_{(i,j-i+1)}$, hence $\operatorname{rk}([i-1,j])=r+1$. Since $\pi^{-1}(j)\in[i,j]$, if $\pi^{-1}(j)=j$, then $j$ is a loop and $\operatorname{rk}([i,j-1])=\operatorname{rk}([i,j])$. If $\pi^{-1}(j)\in[i,j-1]$, then there is a dot in $P_{(i,j-i)}\setminus P_{(i,j-i+1)}$ as well as one dot in $P_{(i,j-i+1)}\setminus P_{(i,j-i)}$, hence $\operatorname{rk}([i,j-1])=\operatorname{rk}([i,j])$. Finally, since $\pi^{-1}(j+1)<i$, the squares in the left-most antidiagonal intersecting $P_{(i,j-i+1)}$ do not contain a dot. Hence, $\operatorname{rk}([i,j+1])$ is given by the number of dots in $P_{(i,j-i+1)}$ plus the number of dots contained in the last row of $P_{(i,j-i+2)}$, i.e. $\operatorname{rk}([i,j+1])=r+1$. It follows that $[i,j]$ satisfies the conditions in eq. 1, i.e., $(r,[i,j])\in\mathcal{E}^{r}$. ∎
@@
Let $\mathcal{P}$ be a positroid and $\mathcal{E}^{r}$ its ranked essential family. For any $[i,j]\subseteq[n]$, we have:
$\operatorname{rk}([i,j])=\min\{r+|[i,j]\setminus I|\mid(r,I)\in\mathcal{E}^{r}\cup\{(0,\emptyset)\}\}.$
@@
Suppose $m_{[i,j]}=|[i,j]|$ and suppose by contradiction that $\operatorname{rk}([i,j])<m_{i,j}=|[i,j]|$. Then there exists a circuit $C\subseteq[i,j]$ (a minimally dependent set). Let $a=\min_{<_{i}}C$ and $b=\max_{<_{i}}C$ and note that $[a,b]$ is such that $\operatorname{rk}([a,b])=\operatorname{rk}([a+1,b])=\operatorname{rk}([a,b-1])$. Then there exists an essential set $(r,I)$ with $r=\operatorname{rk}([a,b])\leq|[a,b]|-1$ and $[a,b]\subseteq I$. Note that $r+|[i,j]\setminus I|\leq|[a,b]|-1+|[i,j]\setminus[a,b]|<|[i,j]=m_{[i,j]}$. Hence, we get a contradiction and $\operatorname{rk}([i,j])=m_{[i,j]}=|[i,j]|$.
Suppose that $m_{[i,j]}=r+|[i,j]\setminus I|$ for some essential set $(r,I)$. Again, suppose by contradiction that $\operatorname{rk}([i,j])<m_{i,j}$. Suppose $[i,j]\setminus I\neq\emptyset$. Then, there exists a circuit $C\subseteq[i,j]$ with $C\not\subseteq I$. Let $a=\min_{<_{i}}C$ and $b=\max_{<_{i}}C$ and note that $[a,b]$ is such that $\operatorname{rk}([a,b])=\operatorname{rk}([a+1,b])=\operatorname{rk}([a,b-1])$ and at least one index between $a$ and $b$ is not in $I$. Then, there exists an essential set $(r^{\prime},J)$ with $[a,b]\cup I\subseteq J$ and $r^{\prime}\leq r+|(J\cap[i,j])\setminus I|-1$. Moreover, we have that:
@@
An essential set $(r,[i,j])\in\mathcal{E}^{r}$ is said to be connected if there exist no pairwise disjoint sets $(r_{1},[i_{1},j_{1}]),\dots,(r_{m},[i_{m},j_{m}])\in\mathcal{E}^{r}$ such that
@@
We denote the set of connected ranked essential sets by $\mathcal{C}\mathcal{E}^{r}$.
@@
rk ([$\textrm{rk }([i,j])=\min\left\{\sum_{a=1}^{m}r_{a}+\biggl{|}[i,j]\setminus\bigcup_{a=1}^{m}[i_{a},j_{a}]\biggr{|}\mid(r_{a},[i_{a},j_{a}])\in\mathcal{C}\mathcal{E}^{r}\cup\{(0,\emptyset)\}\right\}$
@@
Let $\mathcal{P}$ be a positroid and let $\mathcal{E}^{r}$ be its ranked essential family. Then the positroid polytope of $\mathcal{P}$ is given by the facets:
$Q_{\mathcal{P}}=\left\{x\in\mathbb{R}^{n}\mid 0\leq x_{i}\leq 1\;\forall i\in[n],\;\sum_{i\in[n]}x_{i}=k,\;\sum_{\ell\in[i,j]}x_{\ell}\leq r\quad\forall(r,[i,j])\in\mathcal{C}\mathcal{E}^{r}\right\}.$
@@
Let $P_{\mathcal{P}}$ be the associated positroid polytope. Note that $x\in P_{\mathcal{P}}$ satisfies all the inequalities defining $Q_{\mathcal{P}}$, hence $P_{\mathcal{P}}\subseteq Q_{\mathcal{P}}$. Let $x\in Q_{\mathcal{P}}$ and $[i,j]\subseteq[n]$. By Remark 3.11, we have that $\operatorname{rk}([i,j])=\sum_{a=1}^{m}r_{a}+\bigl{|}[i,j]\setminus\bigcup_{a=1}^{m}[i_{a},j_{a}]\bigr{|}$ for some pairwise disjoint intervals $(r_{1},[i_{1},j_{1}]),\dots,(r_{m},[i_{m},j_{m}])\in\mathcal{E}^{r}$. Then

### v3

@@ Abstract
Positroids are a family of matroids introduced by Postnikov in the study of non-negative Grassmannians. Postnikov identified several combinatorial objects in bijections with positroids, among which are bounded affine permutations. On the other hand, the notion of essential sets, introduced for permutations by Fulton, was used by Knutson in the study of the special family of interval rank positroids. We generalize Fulton’s essential sets to bounded affine permutations. The bijection of the latter with positroids, allows the study of the relationship between them. From the point of view of positroids, essential sets are maximally dependent cyclic intervals. We define connected essential sets and prove that they give a facet description of the positroid polytope, as well as equations defining the positroid variety. We define a subset of essential sets, called core, which contains minimal rank conditions to uniquely recover a positroid. We provide an algorithm to retrieve the positroid satisfying the rank conditions in the core or any compatible rank condition on cyclic intervals.
@@ 1 Introduction
Matroids are combinatorial objects introduced to encode the concept of linear dependence arising from vector spaces, graphs, algebraic field extensions, and various other contexts. Positroids form a particularly well-behaved class of matroids. On the combinatorial side, Postnikov has identified several families of objects in bijection with positroids, such as bounded affine permutations, plabic graphs, Grassmann necklaces, and Le diagrams [15]. The realization space of a positroid inside the totally non-negative Grassmannian, called a positroid cell, is homeomorphic to an open ball. Positroid cells fit together to form a CW decomposition of the totally non-negative Grassmannian [16]. Positroids have also been studied in relation to problems coming from quantum physics, as linear images of some positroid cells define a tiling of the amplituhedron, an object encoding the geometry of scattering amplitudes [2, 7, 12].
Here, we introduce a new combinatorial description of positroids via ranked essential sets and explain how to derive information about the positroid’s structure from them. While many combinatorial objects studied in relation to positroids encode their data implicitly, ranked essential sets provide an intuitive understanding of the positroid’s dependencies. These dependencies are succinctly encoded by relying on the cyclic order of the ground set. Additionally, ranked essential sets offer insights into the geometry of the positroid’s realization space within the Grassmannian and the totally non-negative Grassmannian. As an example, the dimension of positroid cells can be explicitly computed in terms of the ranked essential sets.
@@ 2 Preliminaries
We begin this section by recalling the definitions of matroids and positroids. Specifically, we recall two among the various equivalent definitions of a matroid: one through its maximally independent sets, and the other through its closed sets [14].
@@
A matroid $\mathcal{M}$ is a pair $(E,\mathcal{B})$ such that $E$ is a finite set, called ground set, and $\mathcal{B}$ is a collection of subsets of $E$, called bases, satisfying:
@@
We will denote the minimal flat containing a subset $S\subseteq E$, or closure of $S$, by $\langle S\rangle$.
@@
Throughout the paper, without loss of generality, we assume $E=\{1,\dots,n\}=[n]$. The family of cardinality $k$ subsets of $[n]$ is denoted by $\binom{[n]}{k}$. For $i\in\mathbb{Z}$, we denote by $i\mod n$ its representative in $[n]$. For any $i,j\in\mathbb{Z}$, we denote by $[i,j]$ the cyclic interval in $[n]$ from $i\mod n$, to $j\mod n$. For instance, if $n=8$, then $[7,10]=[7,2]=\{7,8,1,2\}$. We use $<$ to denote the normal ordering on $\mathbb{Z}$ and $<_{i}$ to denote the ordering on $[n]$ where $i$ is the smallest element, i.e. $i<_{i}i+1<_{i}\dots<_{i}n<_{i}1<_{i}\dots<_{i}i-1$. We extend this order to $\mathbb{Z}$ by defining, for $j,k\in\mathbb{Z}$, $j<_{i}k$ if and only if $j\mod n<_{i}k\mod n$.
Matroids arise as a generalization of the notion of linear dependency. If a matroid is associated to the dependency behavior of a finite set of vectors in a vector space, it is called realizable. Formally, we have the following definition.
@@
Let $\mathbb{K}$ be a field. A matroid $\mathcal{M}=([n],\mathcal{B})$ is said to be realizable over $\mathbb{K}$ if there exist a positive integer $k$ and a $k\times n$ matrix $M$ with coefficients in $\mathbb{K}$ such that for every $I\in\binom{[n]}{k}$
@@
where $M_{I}$ denotes the square $k\times k$ matrix whose columns are the columns indexed by $I$ in $M$. The matrix $M$ is said to be a realizing matrix for $\mathcal{M}$.
A positroid is a matroid realizable over $\mathbb{R}$ with a positivity condition.
@@
A positroid is a matroid $\mathcal{M}$ which is realizable over $\mathbb{R}$ and such that there exists a realizing matrix $M$ for $\mathcal{M}$ with non-negative maximal minors. In other words, for every $I\in\binom{[n]}{k}$, where $M_{I}$ is as defined in Definition 2.4, we have $\det(M_{I})\geq 0$.
Positroids, along with their combinatorial counterparts, have been subject to extensive research [15]. Moreover, they found applications in many problems coming from different branches of science [17]. One important example is the study of the amplituhedron, an object encoding scattering amplitudes, quantities of interest in theoretical physics [3]. Among the combinatorial objects labeling positroids, bounded affine permutations are of particular interest to us.
@@
Consider a positive integer $n$. A bounded affine permutation of size $n$ is a bijection $\pi:\mathbb{Z}\to\mathbb{Z}$ satisfying the following conditions for all $i\in\mathbb{Z}$:
$\pi(i+n)=\pi(i)+n\quad\text{and}\quad i\leq\pi(i)\leq i+n.$
@@
$\pi(i)=\min\{j\geq i\mid i\in\langle[i+1,\dots,j]\rangle\}\quad\text{ for every }i\in\mathbb{Z}.$
In particular, if $i$ is a loop of the positroid, then $\pi(i)=i$, and if $i$ is a coloop of the positroid, then $\pi(i)=i+n$. Moreover, information on the rank of a positroid on cyclic intervals can be extracted from its associated bounded affine permutation. Indeed, by specializing [11, Theorem 3] to the case of cyclic intervals, we obtain the following result.
@@
Let $\mathcal{M}=([n],\mathcal{B})$ be a matroid. The matroid polytope of $\mathcal{M}$ is
@@
where $e_{B}$ is the indicator vector of $B$, i.e. $e_{B}=\sum_{i\in B}e_{i}$, with $e_{i}$ vectors of the standard basis of $\mathbb{R}^{n}$.
@@ 3.1 Essential sets and bounded affine permutations
Consider a $k\times n$ array of squares in the plane. We will always consider the rows modulo $k$, i.e. like a matrix vertically infinitely repeating. The square at coordinates $(i,j)$ corresponds to the square contained in row $i$ and column $j$ of the array. Let $R_{i}$ be the $i$-th row of the array and $A_{(i,j)}$ be the antidiagonal containing the square $(i,j)$. The sub-antidiagonal $\Delta_{(i,j)}$ of the square $(i,j)$ is the set of squares in the array contained in the antidiagonal starting at the square $(i+1,j-1)$ and ending in the first column, with rows considered modulo $k$. Formally, it is defined as:
@@
Let $\pi$ be a bounded affine permutation of size $n$. We associate to it a dotted array as follows: consider an $n\times(n+1)$ array of squares in the plane. For each $i\in[n]$, place a dot inside the square $(i,\pi(i)-(i-1))$, while all other squares remain white. This is well-defined since $i\leq\pi(i)\leq i+n$. We denote by $D(\pi)$ the set of squares in the dotted array of $\pi$ containing a dot, i.e. $D(\pi)=\{(i,\pi(i)-i+1)\mid i\in[n]\}$.
@@
A white square is called a corner if no other white square shares its top-right corner. The essential family $\mathcal{E}(\pi)$ of $\pi$ is defined as the set of cyclic intervals $[i,j]$ for each white corner $(i,j-i+1)$ in the diagram of $\pi$. Each element in $\mathcal{E}(\pi)$ is referred to as an essential set. Equivalently:
@@
The positroid corresponding to $\pi$ has rank $3$ and can be represented by the point-line configuration in Figure 2 (right). Note that the essential family captures the sets lying on a line and the parallel elements. However, to distinguish between them, we need to know their ranks.
@@
Let $\pi$ be a bounded affine permutation of order $n$. We define the ranked essential family $\mathcal{E}^{\mathrm{r}}(\pi)$ of $\pi$ to be the family
$\mathcal{E}^{\mathrm{r}}(\pi)=\left\{(r,[i,j])\mid[i,j]\in\mathcal{E}(\pi)\cup\{[1,n]\},\;r=\text{ number of dots in }P_{(i,j-i+1)}\right\}.$
@@
Let $\mathcal{U}_{k,n}$ be the uniform matroid of rank $k$ on $n$ elements. The associated bounded affine permutation is defined by $\pi_{k,n}(i)=i+k$, for every $i\in\mathbb{Z}$. It is easy to see that the set of corners of its diagram is empty, hence $\mathcal{E}^{\mathrm{r}}(\pi_{k,n})=\{(k,[1,n])\}$.
@@ 3.2 Essential sets and rank function
The ranked essential family can be interpreted as the rank information of maximally dependent cyclic intervals. Indeed, the essential sets can be constructed from the rank function of the positroid as follows.
@@
Then $\mathcal{E}^{\mathrm{r}}=\mathcal{R}\cup\{(k,[1,n])\}$.
@@
Let $\pi$ be the bounded affine permutation associated to the positroid $\mathcal{P}$. Let $(r,[i,j])\in\mathcal{R}$. Since $\operatorname{rk}([i,j])=\operatorname{rk}([i+1,j])$, we have $\pi(i)\leq j$. Similarly, since $\operatorname{rk}([i-1,j])=\operatorname{rk}([i,j])+1$ and $i-1\not\in\langle i,\dots,j\rangle$, we have $\pi(i-1)>j$. It remains to prove that $\pi^{-1}(j)\leq_{i}j$ and $\pi^{-1}(j+1)>_{i}j+1$. The condition $\operatorname{rk}([i,j])=\operatorname{rk}([i,j-1])$ is equivalent to the fact that the number of dots in $P_{(i,j-i+1)}$ and in $P_{(i,j-i)}$ is the same. Then either $\pi(j)=j$ or $\pi(j)>j$. In the latter case, $P_{(i,j-i+1)}$ contains a dot in the last row, which is not part of $P_{(i,j-i)}$. Hence there must be a dot in $P_{(i,j-i)}\setminus P_{(i,j-i+1)}$, i.e. $\pi^{-1}(j)<_{i}j$. It follows that $\pi^{-1}(j)\leq_{i}j$. Finally, $\operatorname{rk}([i,j])=\operatorname{rk}([i,j+1])-1$ implies that the number of dots in $P_{(i,j-i+2)}$ is one more than the number of dots in $P_{(i,j-i+1)}$. Note that $j+1$ is not a loop, hence $\pi(j+1)>j+1$, and $P_{(i,j-i+2)}$ contains a dot in the row $j+1$, which is not in $P_{(i,j-i+1)}$. For the equality to hold there must be no dot in $P_{(i,j-i+1)}\setminus P_{(i,j-i+2)}$, i.e. $\pi^{-1}(j+1)>_{i}j+1$. It follows that $[i,j]$ is an essential set of $\pi$.
Conversely, let $(r,[i,j])\in\mathcal{E}^{\mathrm{r}}\setminus\{(k,[1,n])\}$. By definition, $r=\operatorname{rk}([i,j])$ is the number of dots in $P_{(i,j-i+1)}$. Moreover, $\operatorname{rk}([i+1,j])$ is equal to the number of dots in $P_{(i+1,j-i)}$, which is equal to the number of dots in $P_{(i,j-i+1)}$ since there is no dot in the first row of this region by definition of essential set. Therefore, $\operatorname{rk}([i+1,j])=r$. In the same way, since $\pi(i-1)>j$, the number of dots in $P_{(i-1,j-i+2)}$ is equal to one more than the dots in $P_{(i,j-i+1)}$, hence $\operatorname{rk}([i-1,j])=r+1$. Since $\pi^{-1}(j)\in[i,j]$, if $\pi^{-1}(j)=j$, then $j$ is a loop and $\operatorname{rk}([i,j-1])=\operatorname{rk}([i,j])$. If $\pi^{-1}(j)\in[i,j-1]$, then there is a dot in $P_{(i,j-i)}\setminus P_{(i,j-i+1)}$ as well as one dot in $P_{(i,j-i+1)}\setminus P_{(i,j-i)}$, hence $\operatorname{rk}([i,j-1])=\operatorname{rk}([i,j])$. Finally, since $\pi^{-1}(j+1)<i$, the squares in the left-most antidiagonal intersecting $P_{(i,j-i+1)}$ do not contain a dot. Hence, $\operatorname{rk}([i,j+1])$ is given by the number of dots in $P_{(i,j-i+1)}$ plus the number of dots contained in the last row of $P_{(i,j-i+2)}$, i.e. $\operatorname{rk}([i,j+1])=r+1$. It follows that $[i,j]$ satisfies the conditions in eq. 1, i.e. $(r,[i,j])\in\mathcal{E}^{\mathrm{r}}$. ∎
@@
Let $\mathcal{P}$ be a positroid and $\mathcal{E}^{\mathrm{r}}$ its ranked essential family. For any $[i,j]\subseteq[n]$, we have:
$\operatorname{rk}([i,j])=\min\{r+|[i,j]\setminus I|\mid(r,I)\in\mathcal{E}^{\mathrm{r}}\cup\{(0,\emptyset)\}\}.$
@@
Suppose $m_{[i,j]}=|[i,j]|$ and suppose for contradiction that $\operatorname{rk}([i,j])<m_{i,j}=|[i,j]|$. Then there exists a circuit $C\subseteq[i,j]$ (a minimally dependent set). Let $a=\min_{<_{i}}C$ and $b=\max_{<_{i}}C$ and note that $[a,b]$ is such that $\operatorname{rk}([a,b])=\operatorname{rk}([a+1,b])=\operatorname{rk}([a,b-1])$. Then there exists an essential set $(r,I)$ with $r=\operatorname{rk}([a,b])\leq|[a,b]|-1$ and $[a,b]\subseteq I$. Note that $r+|[i,j]\setminus I|\leq|[a,b]|-1+|[i,j]\setminus[a,b]|<|[i,j]=m_{[i,j]}$. Hence, we get a contradiction and $\operatorname{rk}([i,j])=m_{[i,j]}=|[i,j]|$.
Suppose that $m_{[i,j]}=r+|[i,j]\setminus I|$ for some essential set $(r,I)$. Again, suppose for contradiction that $\operatorname{rk}([i,j])<m_{i,j}$. Suppose $[i,j]\setminus I\neq\emptyset$. Then, there exists a circuit $C\subseteq[i,j]$ with $C\not\subseteq I$. Let $a=\min_{<_{i}}C$ and $b=\max_{<_{i}}C$ and note that $[a,b]$ is such that $\operatorname{rk}([a,b])=\operatorname{rk}([a+1,b])=\operatorname{rk}([a,b-1])$ and at least one index between $a$ and $b$ is not in $I$. Then, there exists an essential set $(r^{\prime},J)$ with $[a,b]\cup I\subseteq J$ and $r^{\prime}\leq r+|(J\cap[i,j])\setminus I|-1$. Moreover, we have that:
@@
An essential set $(r,[i,j])\in\mathcal{E}^{\mathrm{r}}$ is said to be connected if there exist no pairwise disjoint sets $(r_{1},[i_{1},j_{1}]),\dots,(r_{m},[i_{m},j_{m}])\in\mathcal{E}^{\mathrm{r}}$ such that
@@
We denote the set of connected ranked essential sets by $\mathcal{C}\mathcal{E}^{\mathrm{r}}$.
@@
$\textrm{rk }([i,j])=\min\left\{\sum_{a=1}^{m}r_{a}+\biggl{|}[i,j]\setminus\bigcup_{a=1}^{m}[i_{a},j_{a}]\biggr{|}\text{ such that }\{(r_{a},[i_{a},j_{a}])\}_{a=1,\ldots,m}\subseteq\mathcal{C}\mathcal{E}^{\mathrm{r}}\cup\{(0,\emptyset)\}\right\}$
@@
Let $\mathcal{P}$ be a positroid and let $\mathcal{E}^{\mathrm{r}}$ be its ranked essential family. Then the positroid polytope of $\mathcal{P}$ is given by the facets:
$Q_{\mathcal{P}}=\left\{x\in\mathbb{R}^{n}\mid 0\leq x_{i}\leq 1\;\forall i\in[n],\;\sum_{i\in[n]}x_{i}=k,\;\sum_{\ell\in[i,j]}x_{\ell}\leq r\quad\forall(r,[i,j])\in\mathcal{C}\mathcal{E}^{\mathrm{r}}\right\}.$
@@
Let $P_{\mathcal{P}}$ be the associated positroid polytope. Note that $x\in P_{\mathcal{P}}$ satisfies all the inequalities defining $Q_{\mathcal{P}}$, hence $P_{\mathcal{P}}\subseteq Q_{\mathcal{P}}$. Let $x\in Q_{\mathcal{P}}$ and $[i,j]\subseteq[n]$. By Remark 3.11, we have that $\operatorname{rk}([i,j])=\sum_{a=1}^{m}r_{a}+\bigl{|}[i,j]\setminus\bigcup_{a=1}^{m}[i_{a},j_{a}]\bigr{|}$ for some pairwise disjoint intervals $(r_{1},[i_{1},j_{1}]),\dots,(r_{m},[i_{m},j_{m}])\in\mathcal{E}^{\mathrm{r}}$. Then